Please solve the following by elimination...
3x + 4y = -8
x + 4y = 4
5x + 4y = 22
5x + y = 13
Thank you

Answer by mananth(16946)

3 x + 4 y = -8 .............1
1 x + 4 y = 4 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-3 x -4 y = 8
1 x 4 y = 4
Add the two equations
-2 x = 12
/ -2
x = -6.00
plug value of x in (1)
3 x + 4 y = -8
-18 + 4 y = -8
4 y = -8 + 18
4 y = 10
y = 10/4
y=5/2
----------------
5 x + 4 y = 22 .............1
5 x + 1 y = 13 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 4
-5 x -4 y = -22
20 x 4 y = 52
Add the two equations
15 x = 30
/ 15
x = 2
plug value of x in (1)
5 x + 4 y = 22
10 + 4 y = 22
4 y = 22 -10
4 y = 12
y = 3
m.ananth@hotmail.ca